A student had multiplied a number by 4/3 instead of dividing it by 4/3 and got 70 more than the correct answer. Find the number.


1
Expert's answer
2021-08-19T15:13:41-0400

If the number is x, then we get the next equation:

4/3x = 3/4x + 70,

(4/3 - 3/4)x = 70,

7/12x = 70,

x = 70×12/7 = 120.